OpenStudy (pink.1d):

Compare P(A∩B) with P(A∪B), and explain what each probability means in the context of the situation and data you collected.

OpenStudy (pink.1d):

This is the table

OpenStudy (pink.1d):

Use social media regularly Don't use social media Younger than 18 16 4 Older than 18 11 9

OpenStudy (pink.1d):

a is younger than 18 and b is using social media regularly

OpenStudy (pink.1d):

P(A∩B) = Probability ( younger than 18 & uses social media regularly) = 16 / 40 P(A∪B) = Probability ( younger than 18 OR uses social media regularly) = 31 / 40 is the answer

OpenStudy (pink.1d):

PeerAnswer is better